OpenStudy (anonymous):

A sphere has a radius of 3 and a hemisphere has a radius of 6. Compare the ratio of the volume of the sphere to the volume of the hemisphere.

OpenStudy (anonymous):

volume of the sphere is (4/3)*pi*r^3

OpenStudy (anonymous):

\[(4/3)*\pi*r ^{3}\]

OpenStudy (anonymous):

volume of the hemisphere is \[(2/3)*\pi*r ^{3}\]

OpenStudy (campbell_st):

the hemisphere has a volume \[V = 2/3 \pi 6^3 \] the sphere has a volume \[V = 4/3 \pi 3^3\]

OpenStudy (campbell_st):

to the ratio of the sphere to the hemisphere is \[36\pi : 144 \pi\] just simplify the ratios

OpenStudy (anonymous):

Thanks so much guys. big help